OpenStudy (anonymous):

Find the equation of the tangent to the curve y=2sin^2(x) at x=pi/4

OpenStudy (anonymous):

first find the slope of the tangent using derivatives

OpenStudy (anonymous):

see given y= 2*(sinx)^2 so dy/dx = 2*2(sinx)*cosx at x=pi/4 ,y = 2*(1/sqrt(2)^2) =1 dy/dx =4* 1/(sqrt( 2) * sqrt( 2) ) =2 dy/dx = 2 so we need aline whih has a slope 2 and passes through pi/4,1 so we have eq of reqd tangent y-1= 2*(x- pi/4) y-1= 2*(x-22/28) and so on [using pi=22/7]